# Bayesian Ensembling: Insights from Online Optimization and Empirical Bayes

### v1

@@ 5.4 Online Regression in Non-Stationary Environments
A common issue in the online setting is non-stationarity, possibly from covariate shift or concept drift. To illustrate the effectiveness of OBS in this setting, we apply OBS to the dynamic online ensemble of basis expansions (DOEBE) (Waxman & Djurić, 2024), which uses O-BMA to ensemble several online Gaussian process-based models. The DOEBE uses a linear basis approximation to GPs using random Fourier features (Lázaro-Gredilla et al., 2010; Rahimi & Recht, 2007), and models non-stationary processes using a random walk on these linear parameters, using variance $\sigma_{\text{rw}}^{2}$. This parameter is set to a default value in Waxman & Djurić (2024), but is found to be quite important in performance on several real-world datasets.
We apply OBS instead of O-BMA to ensemble RFF-GPs with $\sigma_{\text{rw}}=10^{-k}$, for $k\in\{0,1,2,3,4\}$. We use the same datasets as Waxman & Djurić (2024) (excluding purely synthetic ones); a summary of datasets and experimental details is in Appendix G. Of note are SARCOS and Kuka #1, both of which are robotics datasets with covariate shift. For SARCOS and Kuka #1 only, we use the smoothed version of EG (Helmbold et al., 1998) with hyperparameters $\eta=10^{-3}$ and $\delta=10^{-2}$; otherwise, EG exhibits severe instabilities.
We include numerical results in Table 1 and figures in Appendix G. We again find OBS to be beneficial. Though EG might be sensitive to outliers and occasionally performs poorly (particularly on SARCOS and Kuka #1), ONS and Soft-Bayes are rather robust and consistently outperform O-BMA. DMA again outperforms O-BMA, and D-ONS performs best on almost every dataset. Interesting future work includes analyzing adaptive regret in non-stationary settings.
@@
Online Baselines O-BMA $-0.59_{\scriptstyle-0.60}^{\scriptstyle-0.58}$ $0.77_{\scriptstyle 0.74}^{\scriptstyle 0.78}$ $0.61_{\scriptstyle 0.52}^{\scriptstyle 0.62}$ $-0.84_{\scriptstyle-0.85}^{\scriptstyle-0.83}$ $0.33_{\scriptstyle 0.32}^{\scriptstyle 0.34}$
DMA $-0.59_{\scriptstyle-0.59}^{\scriptstyle-0.58}$ $0.79_{\scriptstyle 0.77}^{\scriptstyle 0.80}$ $0.69_{\scriptstyle 0.63}^{\scriptstyle 0.70}$ $-0.81_{\scriptstyle-0.81}^{\scriptstyle-0.80}$ $0.33_{\scriptstyle 0.33}^{\scriptstyle 0.34}$
Offline BCRP -$\textbf{-0.57}_{\scriptstyle-0.58}^{\scriptstyle-0.57}-$ $0.80_{\scriptstyle 0.77}^{\scriptstyle 0.81}$ $0.66_{\scriptstyle 0.62}^{\scriptstyle 0.68}$ $-0.81_{\scriptstyle-0.81}^{\scriptstyle-0.81}$ $\textbf{0.35}_{\scriptstyle 0.35}^{\scriptstyle 0.37}$
Online Bayesian Stacking EG $\underline{-0.58}_{\scriptstyle-0.58}^{\scriptstyle-0.57}$ $0.78_{\scriptstyle 0.59}^{\scriptstyle 0.79}$ $0.56_{\scriptstyle 0.45}^{\scriptstyle 0.67}$ $-0.81_{\scriptstyle-0.81}^{\scriptstyle-0.81}$ $0.35_{\scriptstyle 0.30}^{\scriptstyle 0.36}$
Soft-Bayes $-0.58_{\scriptstyle-0.58}^{\scriptstyle-0.58}$ $0.80_{\scriptstyle 0.77}^{\scriptstyle 0.81}$ $\underline{0.70}_{\scriptstyle 0.67}^{\scriptstyle 0.72}$ $-0.81_{\scriptstyle-0.81}^{\scriptstyle-0.81}$ $0.34_{\scriptstyle 0.34}^{\scriptstyle 0.36}$
ONS $-0.58_{\scriptstyle-0.59}^{\scriptstyle-0.58}$ $\underline{0.80}_{\scriptstyle 0.78}^{\scriptstyle 0.82}$ $0.70_{\scriptstyle 0.66}^{\scriptstyle 0.72}$ $\underline{-0.80}_{\scriptstyle-0.81}^{\scriptstyle-0.80}$ $0.34_{\scriptstyle 0.34}^{\scriptstyle 0.35}$
D-ONS $-0.58_{\scriptstyle-0.58}^{\scriptstyle-0.58}$ $\textbf{0.81}_{\scriptstyle 0.79}^{\scriptstyle 0.82}$ $\textbf{0.73}_{\scriptstyle 0.69}^{\scriptstyle 0.74}$$\textbf{-0.80}_{\scriptstyle-0.80}^{\scriptstyle-0.80}-$ $\underline{0.35}_{\scriptstyle 0.35}^{\scriptstyle 0.36}$
@@ 6 Discussion & Conclusions
In this work, we critically study the idea of “Bayesian ensembles,” especially in the online setting. We focused on two strategies for determining the weights in a linear ensemble: Bayesian model averaging (BMA) and Bayesian stacking (BS). In particular, we introduced and discussed online BS (OBS) through the prequential principle. Through careful observation, we established a connection between the OBS problem and the OPS problem, allowing us to leverage the rich literature on online convex optimization. Further, we illustrate how O-BMA optimizes a different loss than OBS leading to a novel empirical Bayesian analysis, and show how O-BMA-based regret bounds can often be adapted to OBS. Finally, we empirically validate our theoretical claims with both synthetic and real datasets, showing that BMA collapses and that the proposed OBS algorithms deliver better performance.
@@
The connection established in this work between OBS and the well-studied problem of portfolio selection bridges optimization and statistics and suggests many interesting future directions. For example, in online or continual learning, one is often concerned with properties under regime changes; a corresponding future direction is to evaluate OBS with algorithms aimed at minimizing dynamic or adaptive regret (Hazan, 2022, Chapter 10). Future statistical investigations include how to initialize ensembles well, and if the same rules of thumb apply as in BMA. It is also interesting to consider connections to non-Bayesian online deep learning methods (Valkanas et al., 2025).
Appendix A OBS Pseudocode
In this appendix, we provide pseudocode for OBS and O-BMA in Algorithm 1. Overall, both methods follow a simple “online inference loop” structure, broken down into a “prediction step” and an “update step.” The difference between OBS and O-BMA is in a single line, underscoring the similarities between these methods.
Algorithm 1 Online Ensemble Update: OBS with EG or OBS with SoftBayes or O-BMA.
2:  for $t=1$ to $T$ do
2:      // Prediction Step
4:     Form ensemble prediction via Equation 1
5:     Output prediction and receive true label $y_{t}$.
5:      // Update Step
6:     Update model weights using OBS w/ EG (10) or OBS w/ SoftBayes (11) or O-BMA (3)
8:  end for
Appendix B More Details on Regret Analysis
@@
One of the main benefits of connecting OBS with OPS is the numerous regret bounds available in the literature. Stating regret bounds for applications of OBS in generality is not possible, as the available regret bounds depend on both the OCO algorithm and the class of models being ensembled. Below, we first state some general bounds reflecting the state of the art in OPS. We then give an example in how these bounds may be applied, showing a regret analysis for OBS applied to an ensemble of approximate Gaussian processes.
B.1 Regret Bounds from Portfolio Selection
@@
In the expert selection problem,
How to apply these bounds to achieve bounds of learning algorithms is an interesting question, and is dependent on the learning algorithm and the type of bound to prove. Below, we provide a simple example that shows OBS may preserve the asymptotic properties of existing regret bounds, even when regret is measured with respect to a single expert rather than a mixture.
B.2 Regret of Online BMA
We now prove Lemma 3.1.
Lemma 3.1.
@@ B.2 Regret of Online BMA
$R_{T}=\sum_{t}\log p_{k^{*}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})-\sum_{t}\log\left(\sum_{k}w_{t,k}p_{k}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1}\right),$ (20)
@@
This follows quickly from the definition of the O-BMA weights Equation 3. Denoting $\ell^{\text{(BMA)}}_{t}=-\sum_{t}\log\left(\sum_{k}w_{t,k}p_{k}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1}\right)$ and $l_{t}^{(k^{*})}=-\log p_{k^{*}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})$, note that
@@
[…]
Taking the logarithm and plugging in the closed form values Equation 2 proves the lemma. ∎
B.3 Example: Online Ensembles of Basis Expansions
A recent example of O-BMA being applied in machine learning is in the ensembling of (approximate) Gaussian processes (GPs). In particular, Lu et al. (2022) proposes ensembling approximate GPs with O-BMA, with Waxman & Djurić (2024) generalizing the algorithm and bounds to more general linear basis expansions. They both prove regret bounds relative to any expert with fixed parameters, a setting where BMA is the appropriate optimizer.
@@ B.3 Example: Online Ensembles of Basis Expansions
Lemma B.1 (Kakade & Ng (2004), Theorem 2.2).
@@
In terms of the time horizon, the bound promised by Lemma B.1 is clearly $\mathcal{O}(\log T)$ in $T$.
@@
We will consider the log-loss of the ensemble at some pre-selected time $T$, denoted $\ell_{T}$, and its regret with respect to the performance of any single expert $k$ with fixed parameter $\boldsymbol{\mathbf{\theta}}_{*}^{(k)}$. Then (a) using O-BMA, the resulting regret is $\mathcal{O}(\log T)$ in $T$; (b) if we further assume that the log-loss is bounded, then the resulting regret remains $\mathcal{O}(\log T)$ using OBS with ONS as the optimizer.
@@
Let $l_{t}^{(k)}$ denote the log-loss of the $k$th expert at time $t$. The proof then proceeds in two steps: first, bounding the loss of the ensemble estimate to any individual expert, then applying Lemma B.1 and combining the bounds.
@@
We proceed by first bounding the ensemble loss to the loss of any individual expert. Beginning with O-BMA, and proceeding identically to proceeding identically to Lemma 3.1, Lu et al. (2022) make the observation that
@@
[…]
@@
where Big O notation is absorbing constants related to $M$ and the maximum gradient norm.
@@
Now, we bound the cumulative loss $\sum_{t=1}^{T}l_{t}^{(k)}$ to the loss using the fixed parameter $\boldsymbol{\mathbf{\theta}}_{*}^{(k)}$. This is achieved by applying Lemma B.1 to the $k$th expert, which results in
@@
Appendix C Experimental Setup

### v2

@@ 5.4 Online Regression in Non-Stationary Environments
A common issue in the online setting is non-stationarity, possibly due to covariate shift or concept drift. To illustrate the effectiveness of OBS in this setting, we apply OBS to the dynamic online ensemble of basis expansions (DOEBE) (Waxman & Djurić, 2024), which uses O-BMA to ensemble several online Gaussian process-based models. The DOEBE employs a linear basis approximation to GPs with random Fourier features (Lázaro-Gredilla et al., 2010; Rahimi & Recht, 2007) and models non-stationary processes by imposing a random walk on these linear parameters, using variance $\sigma_{\text{rw}}^{2}$. This parameter is set to a default value in Waxman & Djurić (2024), but it is found to be quite important for performance on several real-world datasets.
We apply OBS instead of O-BMA to ensemble RFF-GPs with $\sigma_{\text{rw}}=10^{-k}$, for $k\in\{0,1,2,3,4\}$. We use the same datasets as Waxman & Djurić (2024) (excluding purely synthetic ones); a summary of datasets and experimental details is in Appendix˜F. Of note are SARCOS and Kuka #1, both of which are robotics datasets with covariate shift. For SARCOS and Kuka #1 only, we use the smoothed version of EG (Helmbold et al., 1998) with hyperparameters $\eta=10^{-3}$ and $\delta=10^{-2}$; otherwise, EG exhibits severe instabilities.
We include numerical results in Table˜2 and figures in Appendix˜F. We again find OBS to be beneficial. Though EG might be sensitive to outliers and occasionally performs poorly (particularly on SARCOS and Kuka #1), ONS and Soft-Bayes are rather robust and consistently outperform O-BMA. DMA again outperforms O-BMA, and D-ONS performs best on almost every dataset. Interesting future work includes analyzing adaptive regret in non-stationary settings.
@@
Online Baselines O-BMA $0.59_{\scriptstyle-0.60}^{\scriptstyle-0.58}$ $0.77_{\scriptstyle 0.74}^{\scriptstyle 0.78}$ $0.61_{\scriptstyle 0.52}^{\scriptstyle 0.62}$ $0.84_{\scriptstyle-0.85}^{\scriptstyle-0.83}$ $0.33_{\scriptstyle 0.32}^{\scriptstyle 0.34}$
DMA $0.59_{\scriptstyle-0.59}^{\scriptstyle-0.58}$ $0.79_{\scriptstyle 0.77}^{\scriptstyle 0.80}$ $0.69_{\scriptstyle 0.63}^{\scriptstyle 0.70}$ $0.81_{\scriptstyle-0.81}^{\scriptstyle-0.80}$ $0.33_{\scriptstyle 0.33}^{\scriptstyle 0.34}$
Offline BCRP -$\textbf{-0.57}_{\scriptstyle-0.58}^{\scriptstyle-0.57}$ $0.80_{\scriptstyle 0.77}^{\scriptstyle 0.81}$ $0.66_{\scriptstyle 0.62}^{\scriptstyle 0.68}$ $0.81_{\scriptstyle-0.81}^{\scriptstyle-0.81}$ $\textbf{0.35}_{\scriptstyle 0.35}^{\scriptstyle 0.37}$
Online Bayesian Stacking EG $\underline{-0.58}_{\scriptstyle-0.58}^{\scriptstyle-0.57}$ $0.78_{\scriptstyle 0.59}^{\scriptstyle 0.79}$ $0.56_{\scriptstyle 0.45}^{\scriptstyle 0.67}$ $0.81_{\scriptstyle-0.81}^{\scriptstyle-0.81}$ $0.35_{\scriptstyle 0.30}^{\scriptstyle 0.36}$
Soft-Bayes $0.58_{\scriptstyle-0.58}^{\scriptstyle-0.58}$ $0.80_{\scriptstyle 0.77}^{\scriptstyle 0.81}$ $\underline{0.70}_{\scriptstyle 0.67}^{\scriptstyle 0.72}$ $0.81_{\scriptstyle-0.81}^{\scriptstyle-0.81}$ $0.34_{\scriptstyle 0.34}^{\scriptstyle 0.36}$
ONS $0.58_{\scriptstyle-0.59}^{\scriptstyle-0.58}$ $\underline{0.80}_{\scriptstyle 0.78}^{\scriptstyle 0.82}$ $0.70_{\scriptstyle 0.66}^{\scriptstyle 0.72}$ $\underline{-0.80}_{\scriptstyle-0.81}^{\scriptstyle-0.80}$ $0.34_{\scriptstyle 0.34}^{\scriptstyle 0.35}$
D-ONS $0.58_{\scriptstyle-0.58}^{\scriptstyle-0.58}$ $\textbf{0.81}_{\scriptstyle 0.79}^{\scriptstyle 0.82}$ $\textbf{0.73}_{\scriptstyle 0.69}^{\scriptstyle 0.74}$ -$\textbf{-0.80}_{\scriptstyle-0.80}^{\scriptstyle-0.80}$ $\underline{0.35}_{\scriptstyle 0.35}^{\scriptstyle 0.36}$
5.5 Additional Experiments
We include several additional experiments in the appendices. Namely, in Appendix˜G, we perform a sensitivity analysis on the learning rate in OCO algorithms. We come to the general conclusion that the learning rate does not significantly affect the results of our experiments for a set of a priori sensible values.
Similarly, in Table˜5, we include ablation on the forgetting factor hyperparameter in DMA. We find that more aggressive forgetting can help in highly non-stationary environments, but that OBS methods (and, in particular, D-ONS) remain competitive, even without further hyperparameter tuning.
@@ 6 Discussion & Conclusions
In this work, we critically study the idea of “Bayesian ensembles,” especially in the online setting. We focused on two strategies for determining the weights in a linear ensemble: Bayesian model averaging (BMA) and Bayesian stacking (BS). In particular, we introduced and discussed OBS through the prequential principle. Through careful observation, we established a connection between the OBS problem and the OPS problem, allowing us to leverage the rich literature on online convex optimization. Below, we provide some practical guidance and limitations (Section˜6.1), outlook (Section˜6.2), and brief conclusions in Section˜6.3).
6.1 Practical Guidance
When to Prefer OBS.
Our theory and experiments suggest that OBS is preferable in M-open scenarios, i.e., whenever the true data-generating model is not available, which we believe is typically the case in machine learning. When the problem is known to be M-closed, we reiterate previously-known theoretical results that O-BMA is optimal, though OBS, using many different online optimization algorithms, still provides competitive performance. Under anticipated and extreme non-stationarity, DMA with small forgetting factors is usually a viable approach, but otherwise, OBS using OCO algorithms designed for dynamic environments should be preferred.
Choosing Your OCO Algorithm.
The theoretical and practical properties of OBS depend on the OCO algorithm used. Based on our experiments and analysis of the relevant theory, we provide the following recommendations: (1) In scenarios with anticipated non-stationarities, D-ONS should be preferred. (2) In more general scenarios where the anticipated non-stationarity is modest and gradual, guidance is somewhat more subjective. Based on our empirical experiments, we recommend ONS if the computation is not prohibitive, EG if the computation is extremely prohibitive and extreme non-stationarity is not a concern, and Soft-Bayes otherwise; however, we encourage practitioners to experiment in their particular domains.
6.2 Outlook
We view OBS as providing a flexible framework, allowing for exciting research along two different axes.
First, the OBS connection provides an impactful application for new techniques developed in OCO. The connection also motivates a set of potentially interesting assumptions to incorporate into future regret bound analyses. For example, some recent work in OCO has focused on so-called “data-dependent bounds,” which can provide more optimistic regret bounds depending on the statistics of the incoming data (Tsai et al., 2023; Putta & Agrawal, 2025); understanding the statistics of predictive log-likelihoods in certain online Bayesian learning settings could thus be a natural place to develop new bounds. Online learning also serves as a natural place to apply research on switching regret (Pasteris et al., 2024).
Second, OBS is quite a general framework for ensembling Bayesian models. Though we considered a wide range of models and applications in this paper (including variational Bayesian neural networks, GPs via basis expansion approximations, and stochastic volatility models via sequential Monte Carlo), we anticipate many more. For example, our sequential Monte Carlo experiment shows the benefits of OBS over the standard “bank of filters” approach to ensembling particle filters, suggesting novel applications in filtering theory.
It is also interesting to consider connections to other probabilistic machine learning settings, including non-Bayesian online deep learning methods (Valkanas et al., 2025), and applications of distributed or decentralized OCO (Yan et al., 2012; Mateos-Núnez & Cortés, 2014) to decentralized Bayesian inference methods that rely on O-BMA (Llorente et al., 2025).
6.3 Conclusions
Furthermore, we illustrate how O-BMA optimizes a different loss than OBS leading to a novel empirical Bayesian analysis, and show how O-BMA-based regret bounds can often be adapted to OBS. Finally, we empirically validate our theoretical claims with both synthetic and real datasets, showing that BMA collapses and that the proposed OBS algorithms deliver better performance.
@@ 6.3 Conclusions
The connection established in this work between OBS and the well-studied problem of portfolio selection bridges optimization and statistics and suggests many interesting future directions. For example, in online or continual learning, one is often concerned with properties under regime changes; a corresponding future direction is to evaluate OBS with algorithms aimed at minimizing dynamic or adaptive regret (Hazan, 2022, Chapter 10). Future statistical investigations include how to initialize ensembles well and whether the same rules of thumb apply as in BMA.
Acknowledgments
This work was supported by the National Science Foundation under Award 2212506.
Appendix A More Details on Regret Analysis
@@
One of the main benefits of connecting OBS with OPS is the numerous regret bounds available in the literature. Stating regret bounds for applications of OBS in general is not possible, as the available regret bounds depend on both the OCO algorithm and the class of models being ensembled. Below, we first state some general bounds reflecting the state of the art in OPS. We then provide an example of how these bounds may be applied, showing a regret analysis for OBS applied to an ensemble of approximate Gaussian processes.
A.1 Regret Bounds from Portfolio Selection
@@
How to apply these bounds to achieve results for learning algorithms is an interesting question and depends on the learning algorithm and the type of bound to prove. Below, we provide a simple example that shows OBS may preserve the asymptotic properties of existing regret bounds, even when regret is measured with respect to a single expert rather than a mixture.
A.2 Regret of Online BMA
We now prove Proposition˜3.1.
Proposition 3.1.
@@ A.2 Regret of Online BMA
$R_{T}=\sum_{t}\log p_{k^{*}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})-\sum_{t}\log\left(\sum_{k}w_{t,k}p_{k}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})\right),$ (24)
@@
This follows quickly from the definition of the O-BMA weights Equation˜3. Denoting $\ell^{\text{(BMA)}}_{t}=-\sum_{t}\log\left(\sum_{k}w_{t,k}p_{k}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1}\right)$ and $l_{t}^{(k^{*})}=-\log p_{k^{*}}(y_{t}\,|\,\boldsymbol{\mathbf{x}}_{t},\mathcal{D}_{t-1})$, note that
@@
$\exp\left(\sum_{t=1}^{T}\ell^{\text{(BMA)}}_{t}-l_{t}^{(k)}\right)=\frac{\cancel{w_{1}^{(k_{*})}}}{w_{0}^{(k_{*})}}\frac{\cancel{w_{1}^{(k_{*})}}}{\cancel{w_{2}^{(k_{*})}}}\cdots\frac{w_{T}^{(k_{*})}}{\cancel{w_{T-1}^{(k_{*})}}}=\frac{w^{(k_{*})}_{T}}{w^{(k_{*})}_{0}}$.
Taking the logarithm and plugging in the closed form values Equation˜2 proves the proposition. ∎
A.3 Example: Online Ensembles of Basis Expansions
A recent example of O-BMA being applied in machine learning is in the ensembling of (approximate) Gaussian processes (GPs). In particular, Lu et al. (2022) proposes ensembling approximate GPs with O-BMA, with Waxman & Djurić (2024) generalizing the algorithm and bounds to more general linear basis expansions. They both prove regret bounds relative to any expert with fixed parameters, a setting in which BMA is the appropriate optimizer.
@@ A.3 Example: Online Ensembles of Basis Expansions
Lemma A.1 (Kakade & Ng (2004), Theorem 2.2).
@@
In terms of the time horizon, the bound promised by Lemma˜A.1 is clearly $\mathcal{O}(\log T)$ in $T$.
@@
We will consider the log-loss of the ensemble at some pre-selected time $T$, denoted $\ell_{T}$, and its regret with respect to the performance of any single expert $k$ with a fixed parameter $\boldsymbol{\mathbf{\theta}}_{*}^{(k)}$. Then (a) using O-BMA, the resulting regret is $\mathcal{O}(\log T)$ in $T$; (b) if we further assume that the log-loss is bounded, then the resulting regret remains $\mathcal{O}(\log T)$ using OBS with ONS as the optimizer.
@@
Let $l_{t}^{(k)}$ denote the log-loss of the $k$th expert at time $t$. The proof then proceeds in two steps: first, bounding the loss of the ensemble estimate to any individual expert, and then applying Lemma˜A.1 and combining the bounds.
@@
We proceed by first bounding the ensemble loss to the loss of any individual expert. Beginning with O-BMA, and proceeding identically to Proposition˜3.1, Lu et al. (2022) make the observation that
@@
$\exp\left(-\sum_{t=1}^{T}\ell^{\text{(BMA)}}_{t}+l_{t}^{(k)}\right)=\frac{w_{0}^{(k)}}{\cancel{w_{1}^{(k)}}}\frac{\cancel{w_{1}^{(k)}}}{\cancel{w_{2}^{(k)}}}\cdots\frac{\cancel{w_{T-1}^{(k)}}}{w_{T}^{(k)}}=\frac{1}{Mw^{(k)}_{T}}$.
@@
where Big O notation absorbs constants related to $M$ and the maximum gradient norm.
@@
Now, we bound the cumulative loss $\sum_{t=1}^{T}l_{t}^{(k)}$ to the loss using the fixed parameter $\boldsymbol{\mathbf{\theta}}_{*}^{(k)}$. This is achieved by applying Lemma˜A.1 to the $k$th expert, which results in
@@
Appendix B Experimental Setup